Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2014-01-25

Observers: Giorgos Leloudas, Mattias Ergon
Support: Jesper Sollerman, Francesco Taddia, Joel Johansson
TAT on duty: Annalisa de Cia, Maria Teresa Botticella

--------------------------------------------------------------

WEATHER CONDITIONS:
start: wind 5.2 m/s; temp. 10.4 C; humid. 83%; Clear. But few thin cirrus low in horizon.
middle: wind 9.0 m/s; temp. 10.3 C; humid. 82%; Clear.
end: wind 10.9 m/s; temp. 11.6 C; humid. 65%; Clear.

--------------------------------------------------------------

CALIBRATIONS

Evening:
EFOSC bias, arcs


--------------------------------------------------------------

OBSERVATIONS

00:51: Cannot start before humidity remains below 80% for 30 consecutive minutes. It just dropped to 78%.
02:12: For the last 2 hrs the humidity has been fluctuating above and below the 80% limit.
04:50: Nothing has changed. We are still closed.
07:35: Humidity dropped finally. We are opening !


Target     Type      UT      Grism     Slit     Seeing     Comments
LSQ14ii   Class   07:55     g13        1        1.3         Re-trying 2x1800s (20min in twilight). Included host in the slit. z = 0.138. Type still unknown.








Comments